I've done this problem 5 times and I can't figure it out. Please help...
The winner of the track meet had an average speed of 5 meters per second. The second place runner had an average speed of 4.5 meters per second. If the winner finished 2.2 seconds ahead of the second place runner, how long did it take the winner to cross the finish line?
Using this equation:
Rate of Winner x Time of Winner (t) = Rate of 2nd Place Runner x Time of 2nd Place Runner (t+2.2)

Answer by TimothyLamb(4379)

---
s = d/t
t = d/s
---
x = race distance
y = time of winner
---
time equation for winner: using t = d/s
y = x/5
---
time equation for second place runner: using t = d/s
y + 2.2 = x/4.5
y = x/4.5 - 2.2
---
linear system:
y = x/5
y = x/4.5 - 2.2
---
y = y
so the right-hand sides of the two equations above, are equal to each other:
---
x/5 = x/4.5 - 2.2
x/4.5 - x/5 = 2.2
5x/5*4.5 - 4.5x/5*4.5 = 2.2
5x - 4.5x = 2.2*5*4.5
0.5x = 2.2*5*4.5
x = 2.2*5*4.5/0.5
x = 99 m
---
y = x/5
y = 99/5
y = 19.8
---
answer:
winner's time:
y = 19.8 seconds
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php